OpenStudy (anonymous):

y varies directly as the square of x. When x = 5, y = 5. Find y when x = 10.

OpenStudy (lalaly):

y = kx^2 when y=5 and x=5 5=k*25 k=1/5 when x=10 y=1/5 * 10 = 2

OpenStudy (anonymous):

y = 20 y = 10 y = 100 y = 50

OpenStudy (lalaly):

sorry when x=10 y=1/5 * 100 y=20